# From Evaluation to Enhancement: Large Language Models for Zero-Knowledge Proof Code Generation

### v1

@@ 5.4. RQ6: Robust Generalization
To assess whether ZK-Coder sustains strong performance on harder, more recent, and contamination-free tasks, we evaluate on LiveCodeBench (Jain et al., 2024), a recently introduced benchmark designed to be contamination free. We select all tasks labeled as medium difficulty, filter and adapt them using the methodology introduced in Sec. 3.1.3, resulting in a set of 34 tasks (details refer to the Supplementary Material). This offers a more challenging and realistic testbed for ZK program generation.
Table 4. ZK-Coder Code Generation Performance on the Adapted LiveCodeBench Benchmark.
Language ZK-Coder Baseline
Circom 72.25% 2.97%
Language ZK-Coder Baseline
Noir 82.32% 14.72%
Table 4 presents the results. ZK-Coder achieves 72.25% accuracy on Circom and 82.32% on Noir, dramatically outperforming the baseline scores of 2.97% and 14.72%, respectively. These results indicate that even on a harder and contamination-free benchmark, ZK-Coder continues to deliver strong accuracy and robust generalization.
Answer to RQ6: ZK-Coder’s performance on LiveCodeBench confirms that the pipeline generalizes effectively to harder and contamination-free tasks.
@@ 5.5. RQ7: Failure Analysis
Table 5. Distribution of ZK-Coder’s Code Generation Failures.
Language Benchmark Repair Budget Sketch False False Mixed False
Exceed Incorrect Accept Reject Accept & Reject
Circom Humaneval 46.09% 34.78% 0.87% 14.78% 3.48%
LiveCodeBench 36.26% 37.36% 9.89% 15.38% 1.10%
Noir Humaneval 17.33% 53.33% 5.33% 17.33% 6.67%
LiveCodeBench 23.53% 58.82% 0.00% 17.65% 0.00%
Table 5 reports the distribution of ZK-Coder’s failure cases across Circom and Noir on both HumanEval and LiveCodeBench. Each column corresponds to a distinct failure mode. Repair Budget Exceed indicates cases where the repair budget was exhausted without producing any compilable program. Sketch Incorrect denotes that the generated ZKSL sketch failed hidden test cases, suggesting a misunderstanding of the task requirements. False Accept corresponds to programs that erroneously accept invalid witnesses, while False Reject refers to programs that incorrectly reject valid witnesses. Mixed False Accept & Reject captures programs that simultaneously exhibit both types of errors, reflecting severe constraint mis-specification.
The data reveal that the two dominant sources of failure are Repair Budget Exceed and Sketch Incorrect. In both Circom and Noir, these two together account for over 80% of failures on both benchmarks. Repair Budget Exceed often arises when the model is unable to generate any compilable program, which due to gaps in pretraining exposure to ZK-specific constructs. Sketch Incorrect, on the other hand, point to deeper misunderstandings of the natural language description, where the task is not properly captured in the ZKSL representation. The less frequent categories (false accept, false reject, and mixed errors) typically stem from incorrect constraint enforcement when translating a valid sketch into a concrete circuit. These errors indicate local implementation mistakes rather than systemic failures of the pipeline.
Answer to RQ7: In summary, Repair Budget Exceed and Sketch Incorrect constitute the most critical bottlenecks for robust end-to-end generation. The former highlights the need for more ZK-specific training data to improve raw generation capabilities, while the latter underscores the difficulty of grounding natural language descriptions in precise constraints.
@@ 6. Related Work
Emerging research has investigated the use of LLMs for generating code in DSLs, a task made challenging by limited training data and specialized syntax (Joel et al., 2024). To mitigate these difficulties, grammar prompting has been proposed to guide LLMs with partial grammars and thereby improve the well-formedness of generated programs (Wang et al., 2023). RAG techniques further enhance DSL generation by supplying relevant API documentation or code samples (Pimparkhede et al., 2024; Bassamzadeh and Methani, 2024). For DSLs where correctness guarantees are critical, approaches based on formal verification (Councilman et al., 2025) and verified lifting (Bhatia et al., 2024) strengthen reliability by coupling generated code with machine-checkable proofs. Our work advances this line of research by focusing on DSLs for zero-knowledge proofs.
@@
The software engineering community has recently begun to explore tools and methodologies for ZKP development. Numerous researches have proposed to improve the security of ZKPs. Static analysis tools such as Circomspect (Dahlgren, 2022), ZKAP (Wen et al., 2024b), QED2 (Pailoor et al., 2023), and halo2-analyzer (Soureshjani et al., 2023) identify specific vulnerability patterns, while testing tools like MTZK (Xiao et al., 2025) and zkFuzz (Takahashi et al., 2025) use fuzzing to expose security issues. In this work, we focus on reducing the burden of writing ZKP programs by leveraging LLMs’ capabilities, complementing these efforts by helping developers audit and refine their code.
7. Discussion, Limitation and Future Work
ZKP Program Complexity.
Program complexity for ZKPs is measured by the size of arithmetic circuits. It is a critical factor in the efficiency and scalability of ZKP systems, directly determining prover cost and overall performance. To reduce that, a range of compilers have been proposed to optimize circuit construction (Stronati et al., 2024; Aztec Labs, 2025; Ozdemir et al., 2022), while new protocols reduce complexity through advanced arithmetizations with custom gates and lookup arguments (Chen et al., 2023; Gabizon et al., 2019). On the programmer side, several techniques further improve efficiency, such as adopting ZK-friendly hash functions (Grassi et al., 2021), gate packing, and aggregating bits into single field elements. For LLM-based code generation, however, producing efficient circuits is more difficult than generating correct ones. As future work, we plan to survey optimization techniques and guide LLMs in incorporating these strategies.
@@ 7. Discussion, Limitation and Future Work
Recent studies show that LLMs struggle with DSLs and other low-resource languages due to the lack of high-quality training data (Joel et al., 2024), a challenge that is particularly acute in the context of ZKPs (Sheybani et al., 2025). This scarcity also limits the feasibility of fine-tuning approaches. Our work takes a step toward mitigating this limitation by systematizing the generation of program examples in ZK DSLs, which can support targeted training and domain adaptation of LLMs. These examples may be used as a complementary resource for future research in ZK code generation.
Threat to Validity.
Our study on language knowledge relies on multiple-choice QA pairs that were initially generated with LLM assistance. Although some of these questions may contain leading hints or incorrect answers, we address this risk through repeated human validation and cross checking, ensuring that any residual noise has limited influence on the results. The evaluation in ZK-Eval employs carefully designed test cases that check both accepting and rejecting behaviors of generated programs and gadget usages. While test cases cannot fully guarantee correctness, they are constructed to include representative corner cases and to enforce soundness and completeness requirements, thereby reducing the likelihood of false positives. Furthermore, the gadget suite used by ZK-Eval and ZK-Coder consists of 35 representative components selected for their prevalence in ZK libraries and for their coverage of core patterns. This set strikes a balance between complexity and tractability, spanning arithmetic, boolean, comparison, etc., and has already demonstrated clear performance improvements in evaluation. Looking forward, the suite is designed for incremental extension, allowing more advanced and cryptographic composites to be incorporated in future work.
@@ 8. Conclusion
We presented ZK-Eval, the first benchmark for systematically assessing LLM capability in ZK program synthesis, and showed that while models master language knowledge, they struggle with gadget construction and end-to-end synthesis. To address these gaps, we introduced ZK-Coder, a retrieval-augmented framework that guides LLMs from abstract constraint sketches to executable ZK code through constraint-guided retrieval and syntax-semantic repair. We show that ZK-Coder significantly improves the code generation capability of LLMs on both Circom and Noir.
Appendix A Illustrative Workflow Example
We demonstrate the workflow using ZK-Coder to generate a ZKP program that can be used to prove the correctness of a Sudoku assignment in Fig. 13. Given a $4\times 4$ Sudoku grid, the goal is to verify that each row, column, and $2\times 2$ block contains all digits 1144 exactly once.
Figure 13. ZK-Coder’s Illustrative Example Workflow on Proving Sudoku Correctness.
In this illustrative workflow example, ZK-Coder first reads the problem description and formulates it in ZKSL, encoding the necessary constraints on the Sudoku grid. The checker validates the sketch and analyzes it to extract the set of constraint gadgets required. For each gadget, relevant implementation hints are retrieved, and an initial Circom implementation is generated. The compiler then checks this program for syntactic validity, returning diagnostics that guide interactive refinement. Once the code compiles, the model is prompted to generate a small test instance. If the resulting prover behavior deviates from expectations, the model is asked to repair the program. This repair loop continues until either the program passes the LLM-supplied tests or the retry budget is exhausted.
We also present two concrete program examples generated by ZK-Coder for this Sudoku problem, one in Circom (Fig. 14) and one in Noir (Fig. 15). Both programs implement the same verification logic. We annotated the source code with comments to point out key sections, such as input signal declarations, gadget imports or implementations, and composing gadgets with field arithmetic operations to implement the constraints outlined in the ZKSL sketch.
Figure 14. Example Circom Program Generated by ZK-Coder.
Figure 15. Example Noir Program Generated by ZK-Coder.
Appendix B Filtered Tasks in Dataset Curation
As introduced in the main paper, our dataset construction is based on adapting existing programming tasks from HumanEval (Chen et al., 2021) and LiveCodeBench (Jain et al., 2024). To ensure that tasks are suitable for ZK DSLs (no floating-point numbers, no string manipulation, no I/O, deterministic output length, etc.), we filter the tasks using a combination of LLM assisted filtering and manual inspection. We now present the complete set of filtered tasks in Table 6 and Table 7.
Table 6. Filtered tasks from HumanEval.
Filtered Task IDs from HumanEval (Total 68 Tasks)
3 5 8 9 13 24 31 33 35 36 37 39 40 41 42 43 46
49 52 53 55 57 59 60 62 63 65 69 70 72 73 75 76 77
83 85 88 90 94 97 102 106 108 109 110 114 115 116 121 122 123
126 127 128 131 135 138 139 142 145 146 147 150 151 152 155 157 159
Table 7. Filtered tasks from LiveCodeBench.
Filtered Task IDs from LiveCodeBench (Total 34 Tasks)
43 44 62 66 78 94 119 124 129 153 156 172 203 225 232 240 243
250 261 265 280 290 297 305 322 323 336 346 353 367 373 375 384 389
Appendix C Supplementary Results
In this section, we present supplementary results that complement the main findings of the paper but were omitted earlier due to space limitations.
C.1. Error Distribution For Gadget Implementation
Previously, we reported the error distribution averaged across four models. Here, we provide the per-model error distributions. The breakdown by individual model is consistent with the overall trend: OpenAI’s reasoning models outperform open-source models, and the most common error types remain consistent across models.
Figure 16. Error Distribution on GPT-o4-mini.
Figure 17. Error Distribution on GPT-o3.
Figure 18. Error Distribution on DeepSeek-V3.
Figure 19. Error Distribution on Qwen3.
Appendix D Complete Set of Constraint Gadgets
The constraint gadgets used in the main paper during both measurement and augmentation. In measurement, they serve as the targets for evaluating LLMs’ ability to correctly employ or implement gadgets when expressing constraints. In augmentation, they are treated as primitive operators in the sketch language, and in constraint-guided retrieval we prepared knowledge snippets for each gadget to assist LLMs in code generation.
The collection of gadgets is consolidated from three sources: (i) the official Circom standard library (iden3, 2025), (ii) the ZK-kit library (Nico Serrano, [n. d.]), and (iii) the Z3Py API (Research, [n. d.]). The complete list is shown in Table 8. We classify gadgets into five categories: Logical, Bitwise, Comparison, Arithmetic, and Composite, according to their functionality. The typing column specifies input and output types, which are used to disambiguate operators during constraint-guided retrieval. While this set is not exhaustive, it spans a wide range of common constraints in ZK programming. We expect to extend it with additional gadgets in future work.
Table 8. List of constraint gadgets used in measurement and augmentation.
Name Category Typing Source
And Logical Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
And Logical Bool $\times\ldots\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
Or Logical Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
Or Logical Bool $\times\ldots\times$ Bool $\rightarrow$ Bool Circomlib,Z3Py
XOr Logical Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
XOr Logical Bool $\times\ldots\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
Not Logical Bool $\rightarrow$ Bool Circomlib, Z3Py
Or Bitwise Field $\times$ Field $\rightarrow$ Field Z3Py
Or⋄∗ Bitwise Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
And Bitwise Field $\times$ Field $\rightarrow$ Field Z3Py
And⋄∗ Bitwise Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
XOr Bitwise Field $\times$ Field $\rightarrow$ Field Z3Py
XOr⋄∗ Bitwise Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
Equal Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
NotEqual Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
LessThan Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
LessThanOrEqual Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
GreaterThan Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
GreaterThanOrEqual Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
Add Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
Subtract Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
Multiply Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
Divide Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
Modulo Arithmetic Field $\times$ Field $\rightarrow$ Field ZK-Kit, Z3Py
Power Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
FloorDivide Arithmetic Field $\times$ Field $\rightarrow$ Field ZK-Kit, Z3Py
Negate Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
Absolute Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
Inverse Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
Sign Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
Conditional Composite Bool $\times$ Field $\times$ Field $\rightarrow$ Field Circomlib, Z3Py
Conditional Composite Bool $\times$ Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
Sum Composite Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
Product Composite Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
Distinct Composite Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py

### v2

@@ 5.4. RQ6: Robust Generalization
Our adapted HumanEval serves as a controlled benchmark to evaluate the core computation-to-verification shift. To evaluate the robustness of ZK-Coder across varying levels of difficulty and practical contexts, we extend our analysis to two additional datasets: (i) LiveCodeBench (Jain et al., 2024), which focuses on high-complexity algorithmic challenges to assess the limits of the framework’s reasoning; and (ii) a set of Production Case Studies curated from 37 open-source ZKP repositories. To ensure the case studies are representative and avoid selection bias, we employed a reproducible filtering process targeting high-frequency circuit patterns such as Merkle verification, state nullification, etc. Details are provided in the Supplementary Material.
Table 3. Generality Study on ZK-Coder’s Code Generation Performance (Averaged Across Models).
Language LiveCodeBench Production Coding Pattern Case Studies
ZK-Coder Repairing Baseline Simple Baseline ZK-Coder Repairing Baseline Simple Baseline
Circom 44.10% 14.07% 1.72% 90.83% 83.33% 27.50%
Noir 56.93% 33.95% 9.89% 92.09% 89.17% 31.67%
We establish two baselines: the Simple Baseline (direct generation) and the Repairing Baseline (adding the interactive repair loop). Table 3 presents the results. On LiveCodeBench, ZK-Coder achieves a dramatic performance leap, reaching 44.10% on Circom and 56.93% on Noir, significantly outperforming the Repairing Baseline (14.07% and 33.95%). On Production Case Studies, ZK-Coder reaches near-perfect success rates (90.83% on Circom and 92.09% on Noir). While the Repairing Baseline also performs strongly, this high baseline is expected, as these studies focus on standardized, commonly used verification patterns in a controlled context where the high-level logic is well-defined. Nevertheless, ZK-Coder achieves a noticeable gain by bridging the “last mile” of correctness, reducing subtle field-level constraint errors that persist in the baselines.
Answer to RQ6: Results from LiveCodeBench and production case studies confirm that ZK-Coder generalizes effectively across a spectrum of challenges, from high-complexity verification logic to standardized production coding patterns.
@@ 5.5. RQ7: Failure Analysis
Table 4. Distribution of ZK-Coder’s Code Generation Failures (Averaged Across Models).
Language ➀ Repair ➁ Sketched ➂ False Acceptance ➃ False Rejection ➄ Mixed False
Budget Exceed Constraint Incorrect (Under-constrained) (Over-constrained) Acc / Rej
Circom 31.67% 41.56% 1.08% 19.78% 5.92%
Noir 12.60% 73.07% 1.67% 10.89% 1.78%
Table 4 reports the distribution of ZK-Coder’s failure cases. Each column corresponds to a distinct failure mode. Repair Budget Exceed indicates cases where the repair budget was exhausted without producing any compilable program. Sketch Incorrect denotes that the generated constraint sketch failed hidden test cases, suggesting an incorrect specification of the verification constraints. False Acceptance corresponds to programs that erroneously accept invalid witnesses, while False Rejection refers to programs that incorrectly reject valid witnesses. Mixed False Acc / Rej captures programs that simultaneously exhibit both types of errors, reflecting severe constraint mis-specification.
The data reveal that the two dominant sources of failures are ➀ and ➁. In both Circom and Noir, these two together account for over 70% of failures. Failure ➀ often arises when the model is unable to generate any compilable program, which is due to gaps in pretraining exposure to ZK-specific constructs. Failure ➁, on the other hand, reflects errors in correctly modeling the verification constraints from natural language descriptions. The less frequent categories (➂, ➃, ➄) typically stem from incorrect constraint enforcement when translating a valid sketch into a concrete circuit. These errors indicate local implementation mistakes rather than systemic failures of the pipeline.
Answer to RQ7: In summary, Repair Budget Exceed and Sketched Constraint Incorrect constitute the most critical bottlenecks. The former highlights the need for more ZK-specific training data to improve raw generation capabilities, while the latter underscores the difficulty of grounding natural language descriptions in precise verification constraints.
@@ 6. Related Work
Emerging research has investigated the use of LLMs for generating code in DSLs, a task made challenging by limited training data and specialized syntax (Joel et al., 2024; Giagnorio et al., 2025). To mitigate these difficulties, grammar prompting (Wang et al., 2023), RAG techniques (Pimparkhede et al., 2024; Bassamzadeh and Methani, 2024), iterative refinement (Ding et al., 2024), and knowledge transfer methods (Cassano et al., 2024; Mora et al., 2024) have been proposed. While most of these methods (except knowledge transfer methods, which target mainstream computational languages) can be adapted for ZKP code generation, our ablation study (§ 5.3) shows they remain ineffective. ZK-Coder advances this area by designing a system tailored to the constraint paradigm of ZKPs.
@@
The software engineering community has recently begun to explore tools and methodologies for ZKP development, with significant emphasis on security. Static analysis tools such as Circomspect (Dahlgren, 2022), ZKAP (Wen et al., 2024a), QED2 (Pailoor et al., 2023), CCC-Check (Kolozyan et al., 2025), and halo2-analyzer (Soureshjani et al., 2023) identify specific vulnerability patterns, while testing tools (Xiao et al., 2025; Takahashi et al., 2025; Chaliasos et al., 2025) use fuzzing to expose security issues. These tools are invaluable for post-hoc analysis and verification, and can be integrated into our system for further security assurance. Our work complements these efforts by focusing on the generation phase, leveraging LLMs’ capabilities to reduce engineering burden.
7. Discussion and Limitations
ZKP Program Efficiency.
Program efficiency for ZKPs is measured by the size of arithmetic circuits, which directly determines prover cost and overall performance. To reduce circuit size, a range of compilers have been proposed to optimize circuit construction (Stronati et al., 2024; Aztec Labs, 2025; Ozdemir et al., 2022), while new protocols reduce cost through advanced arithmetizations with custom gates and lookup arguments (Chen et al., 2023; Gabizon et al., 2019). On the programmer side, several techniques further improve efficiency, such as adopting ZK-friendly hash functions (Grassi et al., 2021), gate packing, and aggregating bits into single field elements. For LLM-based code generation, however, producing efficient circuits is more challenging than generating correct ones. As future work, we plan to survey optimization techniques and guide LLMs on this.
@@ 7. Discussion and Limitations
Recent studies show that LLMs struggle with DSLs and other low-resource languages due to the lack of high-quality training data (Joel et al., 2024), a challenge that is particularly acute in the context of ZKPs (Sheybani et al., 2025). This scarcity also limits the feasibility of fine-tuning approaches. Our work takes a step toward mitigating this limitation by systematizing the generation of program examples in ZK DSLs, which can support targeted training and domain adaptation of LLMs. These examples can serve as a complementary resource for future research in ZK code generation.
Real-World Representativeness.
We examine the representativeness and scope of our evaluation along two complementary dimensions. First, from a structural standpoint, the programs generated in § 5.4 exhibit an average cyclomatic complexity of 22.81, which exceeds the average observed in production-grade open-source ZK repositories (16.8 for Noir and 17.5 for Circom). This indicates that our benchmarks not only mirror but often surpass the structural depth of real-world ZK logic. Second, while our tasks are designed to isolate the core challenge of constraint specification, our evaluation is not “API-blind.” We recognize that LLM capabilities in standard and cryptographic API usage are already well established in prior work (Qin et al., 2023; Zhuo et al., 2024; Wang et al., 2025b; Zhu et al., 2025). Consequently, our test suite focuses on verifying that these APIs are integrated correctly within the circuit logic, covering a considerable number of real-world in-circuit invocations such as cryptographic primitives and data-processing utilities. This dual focus ensures that ZK-Coder is evaluated against the rigorous logical demands of production-level ZK development.
Threats to Validity.
We identify these potential threats. ➀ Our language-knowledge study uses multiple-choice QA items that were initially generated with LLM assistance. To mitigate bias and errors, we performed iterative human validation and cross-checking; while some residual noise may remain, we expect it to have limited impact on the reported trends. ➁ The ZK-Eval assessment uses hand-crafted test cases that exercise both accepting and rejecting behaviors of generated programs and primitive usages. Although tests cannot prove correctness, they include representative corner cases and explicit checks for soundness and completeness, which substantially reduce the risk of false positives. The high mutation detection rate reported in § 3.1.3, achieved through oracle-guided mutation testing, further supports this claim by demonstrating strong sensitivity to subtle constraint specification errors. ➂ The reliability of our automated evaluation depends on test-case coverage. We mitigate this through extensive cross-validation and edge-case injection, achieving a high mutation score. While this is not a formal proof of logical equivalence, it provides strong empirical assurance of correctness. ➃ The choice of prompting strategy may influence LLM performance. We employ standard zero-shot and few-shot prompting techniques without extensive prompt engineering or fine-tuning. To ensure experimental integrity, we verified that no test cases or ground-truth solutions were included in the prompts, preventing information leakage that could artificially inflate success rates. ➄ Security is critical for ZKPs. While integration with auditing tools (Pailoor et al., 2023; Kolozyan et al., 2025) could further enhance assurance, code generation from natural language specifications may face inherent ambiguity and underspecification. Users must exercise caution and validate generated code before deployment in security-critical application.
@@ 8. Conclusion
We presented ZK-Eval, the first benchmark for assessing LLM capability in ZK program generation, revealing that models struggle with composing algebraic primitives for verification constraints. To address this, we introduce ZK-Coder, which guides LLMs from constraint sketches to executable code through sketch-guided retrieval and iterative refinement, significantly improving code generation for both Circom and Noir.